OpenStudy (anonymous):

f(x) varies directly with x and y=-2 when x=8 A.find the constant of variation k B.write the direct variation equation C.find the value of f(x) when x=-4 D.find the value of x when f(x)=4 @ranga

OpenStudy (ranga):

This is similar to the problem we did earlier. y = k * x. when x = 8, y = -2. Solve for k.

OpenStudy (anonymous):

-2=k*8

OpenStudy (ranga):

yes. so what is k?

OpenStudy (anonymous):

then it turns into k=-2*8 right?

OpenStudy (ranga):

no. to solve for k you need to isolate k which can be done by dividing both sides by 8 and simplifying.

OpenStudy (anonymous):

-0.25=k?

OpenStudy (ranga):

Leave it as a fraction instead of decimals.

OpenStudy (anonymous):

so -1/4?

OpenStudy (ranga):

yes. a) k = -1/4

OpenStudy (anonymous):

ok part b now.

OpenStudy (anonymous):

y=k*x?

OpenStudy (ranga):

Yes. But you already solved for k and so put k = -1/4

OpenStudy (anonymous):

but the part b says I have to write the direct variation equation that is the equation right?

OpenStudy (ranga):

y is same as f(x). So you can write it as f(x) = -1/4x (I use * so people don't think it is 1 divide by 4x. But when answering you don't have to use *). I will use the math editor to show the answer properly:

OpenStudy (ranga):

\[f(x) = -\frac{ 1 }{ 4 }x\]

OpenStudy (anonymous):

oh ok makes better sense ready to tackle part c?

OpenStudy (ranga):

C. Put x = -4 in the equation we got for B and find f(x).

OpenStudy (anonymous):

f(4)=-1/4x?

OpenStudy (ranga):

We need to find f(-4) f(x) = -1/4 * x f(-4) = -1/4 * (-4) = 1

OpenStudy (ranga):

For D. Put f(x) = 4 into the equation and solve for x f(x) = -1/4 * x 4 = -1/4 * x multiply both sides by 4 16 = -x x = -16

OpenStudy (anonymous):

Ok thanks I lost my internet for a little bit but thanks for everything I greatly appreciate it! :)

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

you are welcome.